OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and that passes thru the givn point y=9/5x-5;(10,5)

OpenStudy (aroub):

parallel=same slope y=mx+b , m=9/5, (10,5) 5=9/5(10)+b 5=18+b -18+5=b -13=b y=9/5x-13